A copper pipe is 24 m long at 0 oC. If the temperature is increased to 52 oC, by how much (in mm) does its length change?
Linear expansion coefficients α for copper: 17*10-6/oC
Enter a number with two digits behind the decimal point.
Answer #1

∆L (lengt change) = L× × ∆T

Thus , ∆L = 24× 17× 10-6×52 = 21216× 10-6 m = 21.216 mm.